Circular No. 6387

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

EMISSION OBJECT IN CYGNUS E. S. Parsamian, Byurakan Astrophysical Observatory and Instituto

de Astronomía, Universidad Nacional Autonoma de México, reports: “During long-slit (3′′×360′′) spectral observations (range 450–700 nm, resolution 0.55 nm/pixel) of V1478 Cyg on the 2.12-m telescope (+ Landessternwarte Faint Object Spectograph and Camera) at Guillermo Haro Observatory, Cananea, on 1995 July 4.23, 5.37, 6.39, 7.30, 9.21, and Sept. 18.35 UT, I discovered a high-velocity object 33′′ (± 2′′) south of V1478 Cyg = MWC 349 (cf. Hamann and Simon 1988, Ap.J. 327, 876). The object is located at α = 20^{h}32^{m}45^{s}.63, δ = +40^{o}39′03′′.6 (equinox 2000.0) and has an emissionline spectrum and faint red continuum similar to V1478 Cyg. A few identified lines (Hα, He II, He I) show a blueshift of ∼ 1000–2000 km/s (possibly variable from night to night) with respect to the same lines in V1478 Cyg (whose V_{LSR} = +8 km/s). Hα seems to be double-peaked with peak separation ∼ 800 km/s. The source of the blueshifted line is tentatively identified with a weak object seen on the red Palomar Sky Survey plate, 32′′ south of V1478 Cyg. As the object lies close to the axis of outflow from V1478 Cyg, perhaps there is a connection. The relative fluxes in 10^{−13} erg cm^{−2} s^{−1} for the Hα lines (using the star Kopff 27 as a standard) of the new object are: 1995 July 4.227, 3.41; 4.246, 3.76; 5.389, 1.30; 5.399, 93.3.”

4U 1728–34 T. Strohmayer, W. Zhang, A. Smale, and C. Day, Universities Space

Research Association (USRA) and the Laboratory for High Energy Astrophysics, Goddard Space Flight Center (LHEA/GSFC); J. Swank, LHEA/ GSFC; L. Titarchuk, George Mason University and LHEA/GSFC; and U. Lee, Tohoku University, report: “Of the seven bursts that we have observed from 4U 1728–34 during a recent campaign with RXTE, five show oscillations with a frequency of 363 Hz. In the first few seconds of two of the seven bursts, the oscillations do not appear to be coherent. The oscillation frequency drifts by as much as 0.5 Hz. During the other three bursts, the coherence of the oscillations is only limited by the durations of the bursts. The fractional rms of the oscillations ranges from 1.5 to 5 percent. We have also found in the same data set two simultaneously-present kHz quasiperiodic oscillations (QPOs), one of which has been reported earlier (IAUC 6320). The centroid frequencies of the two QPOs change with intensity and time, but their difference appears to be always near 363 Hz. These observations are consistent with a neutron star spin period of 2.75 ms.”

1996 April 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green